# Robust variational neural posterior estimation for simulation-based inference

### v1

@@ 4 Experiments and results
We evaluate our method and test the main claim of our paper: that RVNP and its tuning variant can recover robust amortised posterior inference. In the experiments section, we test four variants of our algorithm
@@
  • RVNP-Global: the standard RVNP algorithm with no final tuning step, using the global error model given in Equation 13.
@@
  • RVNP-Global-T: the tuned version of RVNP-Global.
In each experiment, we only test the tuned algorithm for whichever algorithm performed better, RVNP or RVNP-Global. The architecture and training procedure for each task are described in the Appendix.
Benchmarking Algorithms. In these tasks, we benchmark RVNP against noisy neural posterior estimation (NNPE) from Ward et al. (2022) due to the similar performance of NNPE with RNPE therein and NNPE’s amortisation ability. We do not benchmark against the MMD or consistency loss methods (Elsemüller et al. 2025; Mishra et al. 2025) due to their dynamic adaptation of the neural statistics. We also benchmark against vanilla NPE.
Summary of Results. We find that RVNP and RVNP-T can recover robust posterior inference in an amortised manner across a range of different tasks, including cases where a significant number of points have ID data. Furthermore, we test RVNP and RVNP-T using low-resolution spectra cross-matched with high-resolution spectroscopic data, and we validate our posterior recovery against classically derived results. For each task, except the Spectra task, we test our method for $N_{\rm obs}\in\{1,10,100,1000,10000\}$ points and evaluate the effectiveness of our methods. We note that in the absence of misspecification, RVNP collapses towards NPE and is well-specified and calibrated.
RVNP and RVNP-T can recover robust posterior inference without adopting priors over misspecification or introducing ad hoc hyperparameters that must be tuned. We display examples of posterior inference in the $N_{\rm obs}=1000$ case for RVNP against NPE for each of the tasks for a single example in Figure 2. We argue that the success of RVNP, the interpretability of the error correction model, and the lack of hyperparameters to be tuned (other than the optimiser hyperparameters) point to a significant contribution to robust amortised SBI. In all tasks, our algorithm is overconfident for $N_{\rm obs}=1$. This is not surprising, as without any other information, our model collapses to the NPE solution. With stronger priors over the misspecification, we could make this step more robust, but this defeats the purpose of allowing the data to drive the error model.
@@ 4.1 Metrics for assessing misspecification
We consider three main metrics to assess the robustness of the inference. Let $\bm{\theta}^{*}$ be the true value of the parameter. Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter over the dataset to assess the performance of RVNP. LPP has been extensively used in SBI literature (Papamakarios & Murray 2016; Hermans et al. 2020; Ward et al. 2022;Kelly et al. 2024; Wehenkel et al. 2025).
Given a credible level $\gamma$, let $\rm{HDR}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)$ represent the $1-\gamma$ highest posterior density region of the posterior $p(\bm{\theta}|\,\bm{x}_{\rm obs})$. The expected posterior coverage (EPC) at a given confidence level over a test set is given by
$\rm{EPC}(\gamma):=\mathbb{E}_{\bm{\theta}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta^{*}\in\rm{HDR}_{p(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)\}],$ (16)
where $\mathbf{1}$ is the indicator function. EPC is a commonly used metric to assess robustness and calibration of posterior distributions, particularly when looking at single observation situations. When comparing posterior calibration across a range of amortised datasets, we compute the average expected posterior coverage AEPC
@@
which represents the average calibration across the test set. Finally, we also compute the normalised (root) mean squared error (NMSE):
$\text{NMSE}=\frac{1}{N_{\rm obs}}\sum_{j=1}^{N_{\rm obs}}\frac{\sqrt{\frac{1}{S}\sum_{s=1}^{N_{\rm samples}}\left(\theta^{*}_{j}-\theta^{(s)}_{j}\right)^{2}}}{(\max(\theta_{\rm{prior}})-\min(\theta_{\rm{prior}}))},$ (18)
which evaluates the normalised accuracy of the posterior prediction to the truth relative to the prior width.
@@ 4.2 Task A: CS
We reproduce the cancer and stromal cell development benchmark task from Ward et al. (2022). The simulator models the development of cancer and stromal cells in 2D space based on the locality of a cell relative to unobserved parents. This is emulated conditional on three Poisson rate parameters $\bm{\theta}=(\lambda_{c},\lambda_{p},\lambda_{d})$. The total number of cells $N^{c}$, number of unobserved parents $N^{p}$, and the number of daughter cells for each parent $N_{i}^{d}$ are sampled as $N^{c}\sim\rm{Poisson}(\lambda_{c})$, $N^{p}\sim\rm{Poisson}(\lambda_{p})$, and $N^{d}_{i}\sim\rm{Poisson}(\lambda_{d})$ for $i=1,...,N^{p}$. Cell locations $\{c_{i}\}_{i=1}^{N_{c}}$ and disease origin points $\{p_{i}\}_{i=1}^{N_{p}}$ were sampled uniformly across the 2D domain using homogeneous spatial point processes. For each origin point $p_{i}$, $r_{i}$ is the Euclidean distance to its $N_{d}^{i}$-th nearest cell. Cells falling within or on the boundary of this radius from $p_{i}$ are designated as cancerous. Distance-based summary measures were estimated by randomly sampling 50 stromal cells. The summary statistics are as follows: N Cancer and N Stromal, the number of cancer and stromal cells, respectively; and (Mean Min Dist) and (Max Min Dist), the mean and maximum distance from the stromal cells to their nearest cancer cell, respectively. The Numba just-in-time implementation of this task was taken directly from the data products of Ward et al. (2022).
@@
Results. We display the results for the CS task in Figure 3 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 7. In this task, the misspecification is minimal, and all versions of our algorithm can recover robust posterior inference beginning at the $N_{\rm obs}=10$ point. We find that all variants of RVNP perform similarly to NNPE in this task. This is not surprising, as the misspecification can be described by inflating the covariance along two specific axes. The RVNP algorithms slightly outperform NNPE in NMSE.
Figure 3: Results for the CS task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
@@ 4.3 Task B: SIR
We include the misspecified susceptible-infected-recovered (SIR) task from Ward et al. (2022), which takes the stochastic model of epidemic spread modelled conditional on $\bm{\theta}=(\beta,\gamma)$, the time-varying infection rate and the recovery rate, respectively. The SIR model emulates ideal disease transmission dynamics from the susceptible ($s$), infected ($i$), and recovered ($r$) parameters as
@@
Ward et al. (2022) employs a stochastic extension by using time-dependent transmission dynamics through a variable infection rate $\tilde{\beta}_{t}$, accounting for external factors such as policy interventions or pathogen mutations. This stochastic process is characterized using the basic reproduction number $R_{0t}=\frac{\tilde{\beta}_{t}}{\gamma}$, which follows the mean-reverting stochastic differential equation:
@@
where $\eta$ controls the mean reversion strength of $R_{0t}$ toward the equilibrium value $\frac{\beta}{\gamma}$, $\sigma$ represents the volatility parameter, and $W_{t}$ denotes standard Brownian motion. $\eta=0.05$ and $\sigma=0.05$ are fixed and the goal is to infer the parameters $\beta$ and $\gamma$. The Julia code to sample from this process was taken directly from the data products of Ward et al. (2022). The summary statistics produced in this task are the mean, median and maximum number of infections, the day of the maximum number of infections, and the day at which half of the total number of infections was reached, and the mean autocorrelation of infections with lag 1.
Figure 4: Results for the SIR task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
@@
Results. We display the results for the SIR task in Figure 4 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 8. In this task, the misspecification is extreme but occurs only along the final axis, and all versions of our algorithm can recover robust posterior inference beginning at the $N_{\rm obs}=10$ point. As the misspecification occurs along a given axis, we expect NNPE to perform well in this task. We find that each of our algorithms witnesses slight performance gains in each of the metrics as the number of observations increases.
@@ 4.4 Task C: Pendulum Task
We describe a stochastic pendulum simulator that, given $\bm{\theta}:=[\omega_{0},A]$, samples the horizontal position of a frictionless pendulum at 200 time points evenly sampled every 0.05 seconds $\bm{x}_{\rm sim}=(f(t_{0}),...,f(t_{200}))$ where $f(t)=A\cos(\omega_{0}t+\phi)$ for $\phi\sim U(-\pi,\pi)$. In this task, $\omega_{0}$ and $A$ denote the fundamental frequency and amplitude, respectively, of the frictionless pendulum. $\phi$ is a stochastic phase shift. This task was inspired by the task from Wehenkel et al. (2025). However, it differs significantly and was adjusted to test our claim that increasing the number of posterior observations will better constrain the parameters.
@@
Misspecification We synthesise a time calibration error in the instrumentation that causes the instrument to take 200 measurements every 0.075 seconds instead of the simulated 0.05 seconds. There is a significant probability that the misspecified point will appear ID due to the mild misspecification.
@@
Neural Statistic Estimation In this example, each of the data points is a single observation. The neural statistic estimator is an embedding of the full pendulum time series into a lower-dimensional representation of the data. Following Chen et al. (2021), we use the Shannon-Jensen InfoMax objective (Hjelm et al., 2019) to target sufficient neural statistics, $\iota_{\bm{\omega}}$, to encode the high-dimensional data. This objective function maximises the mutual information between $\bm{x}_{\rm sim}$ and $\bm{\theta}$ using a discriminator network.
@@
Figure 5: Results for the Pendulum task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
@@
Results. We display the results for the Pendulum task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 9. In this task, the misspecification is geometrically significantly more complex. Moreover, due to the nature of the misspecification, most of the points will appear ID. Due to the complexity of the misspecification, NNPE struggles significantly. However, RVNP and RVNP-Global recover robust posterior inference for this task. In particular, RVNP is very well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range in parameter space. Due to the complexity of the error model, the tuned variant, RVNP-T, does not see the same degree of performance increase. However, it does perform better than both NPE and NNPE.

### v2

@@ 4 Experiments and results
We test the main claim of our paper: that RVNP and its tuned variant can recover robust amortised posterior inference. In each task, we chose a fixed simulation budget of $N_{\rm sim}=100,000$ and retained $10\%$ for validation. For the three synthetic tasks, we test our methods for $N_{\rm obs}\in\{1,10,100,1000,10000\}$ points. For the task using real astronomical data, we test our methods for $N_{\rm obs}\in\{1,10,100,1000\}$ due to limitations on the available data. Furthermore, we implemented sample-importance-resampling with $5000$ particles and $100$ samples per particle. In the experiments section, we test three variants of our algorithm
@@
  • RVNP-G: the standard RVNP algorithm with no final tuning step, using the global error model given in Equation 15.
@@
For the RVNP and the RVNP-G algorithm, we evaluate the metrics with and without sample-importance-resampling. For both the variational posterior and the simulator proxy, we used neural spline flows (NeuralSplineFlows), and the hyperparameters were kept consistent among the tasks. An overview of the architecture and training procedure for each task is described in the appendices.
Benchmarking Algorithms. In these tasks, we benchmark RVNP against noisy neural posterior estimation (NNPE) from ward2022robust because NNPE is a robust amortised Bayesian inference algorithm. Furthermore, NNPE exhibits similar performance relative to their robust neural posterior estimation algorithm, which is not suited to amortised inference. We do not benchmark against the unsupervised domain adaptation or consistency loss methods for amortised Bayesian inference (elsemüller2025doesunsuperviseddomainadaptation; mishra2025robust) due to their dynamic adaptation of the embedding space. We also benchmark against standard NPE.
Summary of Results. We found that RVNP and RVNP-T can recover robust posterior inference in an amortised manner across a range of different tasks, including cases where a significant number of points have ID data. RVNP-T does not increase the robustness of the posterior across any task, and should be avoided unless the practitioner believes the error model has been inferred globally and wants to infer posteriors on streaming data in a fast manner. Furthermore, we found that sample-importance-resampling from the variational posterior increases the posterior performance slightly.
Our methods performed well against the benchmark algorithms except for the $N_{\rm obs}=1$ case, where we find that our algorithm is always overconfident to a certain degree. This is not surprising, as RVNP is highly over-parametrised without enough data to learn the error model reliably. With stronger priors on misspecification and reducing the flexibility of the error model, we could make this step more robust. However, this defeats the purpose of allowing the data to drive the error model. Furthermore, we do not attempt to address robustness in the non-amortised inference problem, but include it for completeness.
We show examples of posterior inference in the $N_{\rm obs}=1000$ case for RVNP against NPE for each of the tasks for a single example in Figure 2 We argue that the success of RVNP, the interpretability of the error correction model, and the lack of misspecification-dependent hyperparameters to set, amount a significant contribution to robust amortised SBI.
@@ 4.1 Metrics for assessing misspecification
We consider four metrics to assess the robustness of the inference. Let $\bm{\theta}^{*}$ be the true value of the parameter. Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter for all points in the dataset. LPP has been extensively used in SBI literature to assess performance, and is a useful metric for assessing the ability to predict the truth from observation (2016arXiv160506376P; hermans2020likelihoodfreemcmcamortisedapproximate; ward2022robust;kelly2024misspecificationrobust; wehenkel2025addressing).
Given a credible level $\gamma$, let $\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)$ represent the $1-\gamma$ highest posterior density region of the posterior estimator $q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})$. The expected posterior coverage (EPC) at a given confidence level over a test set is given by
$\rm{EPC}(\gamma):=\mathbb{E}_{\bm{\theta}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta^{*}\in\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}|\,\bm{x}_{\rm obs})}(1-\gamma)\}],$ (16)
where $\mathbf{1}$ is the indicator function. EPC is a commonly used metric to assess robustness and calibration of posterior distributions (kelly2025simulation), particularly when looking at single observation situations. When comparing posterior calibration across a range of amortised datasets, we compute the average expected posterior coverage (AEPC)
@@
which represents the average calibration across the test set. The AEPC metric is evaluated using the density approach of hyndman1996hdr for calculating highest posterior density intervals .
We also compute the average expected marginal posterior coverage (AEMPC) for each parameter $\bm{\theta}_{i}$, which is common when we only obtain samples of the posterior distribution (wehenkel2025addressing). The AMEC can be implemented in a more stable fashion than the posterior density approach but can oversimplify the complex regions of the posterior. It is calculated by first estimating
$\rm{EMPC}(\gamma)_{i}:=\mathbb{E}_{\bm{\theta}_{i}^{*},\bm{x}_{\rm obs}\sim T}[\mathbf{1}\{\theta_{i}^{*}\in\rm{HDR}_{q_{\bm{\phi}}(\bm{\theta}_{i}|\,\bm{x}_{\rm obs})}(1-\gamma)\}].$ (18)
The AEMPC is defined as
$\alpha\,\rm{(marginal)}:=\frac{1}{m}\sum_{i=1}^{m}\int_{0}^{1}[\rm{EMPC}(\gamma)_{i}-\gamma]d\gamma,$ (19)
which is necessary for comparing the coverage of the posterior when using sample-importance-resampling, as we only retrieve the unnormalised posterior associated with sample-importance-resampling distribution. We did not wish to rely on multi-dimensional density estimation methods to evaluate the posterior coverage from samples from the joint posterior, as we found that it depended on the density estimation algorithm.
Finally, we also evaluated the normalised root mean squared error (NRMSE):
$\text{NRMSE}=\frac{1}{N_{\rm obs}}\sum_{j=1}^{N_{\rm obs}}\frac{\sqrt{\frac{1}{S}\sum_{s=1}^{N_{\rm samples}}\left(\theta^{*}_{j}-\theta^{(s)}_{j}\right)^{2}}}{\rm{Std}(\theta_{\rm{prior}})},$ (20)
which evaluates the normalised accuracy of the posterior prediction to the truth relative to the standard deviation of the prior training samples in each dimension.
4.1.1 Description of Plots
We evaluated the metrics on each task and displayed the results in Figures 3, 4, 5, and 6. The purple dashed line corresponds to NNPE, and the orange dashed line corresponds to NPE. The blue points are from RVNP with the darker blue indicating sample-importance-resampling. The green points are from RVNP-G and the darker green indicates sample-importance-resampling. The red points corresponds to the tuned variant RVNP. In Appendix A.2, we show how RVNP performs on well-specified tasks.
The top-left plot in each image shows the AEPC ($\alpha$) as a function of the number of observations, with the light green hue showing the desired target region. A value of 0 in this metric indicates well-calibrated. Any values less than this implies overconfidence, and positive results imply posterior under-confidence. Under-confidence is more desirable than overconfidence. The values of NPE and NNPE are independent of the number of observations because they are not tuned dynamically. The top-right plot in each image shows the LPP as a function of the number of observations, with the error bars indicating the standard deviation. The purple hue indicates the standard deviation of LLP from NNPE, and the orange hue indicates the standard deviation of LLP from NPE. A higher value of LPP is better. The bottom-left plot in each image shows the AMEPC ($\alpha$ marginal) as a function of the number of observations, with the light green hue showing the desired target region. Similarly to AEPC, a value of 0 in this metric indicates well-calibrated. In the bottom-right plot, we show the NRMSE as a function of the number of observations, where lower values are better.
@@ 4.2 Task A: CS
We reproduce the cancer and stromal cell development benchmark task from ward2022robust. The simulator models the development of cancer and stromal cells in 2D space based on the locality of a cell relative to unobserved parents. This is emulated conditional on three Poisson rate parameters $\bm{\theta}=(\lambda_{c},\lambda_{p},\lambda_{d})$. The total number of cells $N^{c}$, number of unobserved parents $N^{p}$, and the number of daughter cells for each parent $N_{i}^{d}$ are sampled as $N^{c}\sim\rm{Poisson}(\lambda_{c})$, $N^{p}\sim\rm{Poisson}(\lambda_{p})$, and $N^{d}_{i}\sim\rm{Poisson}(\lambda_{d})$ for $i=1,...,N^{p}$. Cell locations $\{c_{i}\}_{i=1}^{N_{c}}$ and disease origin points $\{p_{i}\}_{i=1}^{N_{p}}$ were sampled uniformly across the 2D domain using homogeneous spatial point processes. For each origin point $p_{i}$, $r_{i}$ is the Euclidean distance to its $N_{d}^{i}$-th nearest cell. Cells falling within or on the boundary of this radius from $p_{i}$ are designated as cancerous. Distance-based summary measures were estimated by randomly sampling 50 stromal cells. The summary statistics are as follows: N Cancer and N Stromal, the number of cancer and stromal cells, respectively; and (Mean Min Dist) and (Max Min Dist), the mean and maximum distance from the stromal cells to their nearest cancer cell, respectively. The Numba just-in-time implementation of this task was taken directly from the data products of ward2022robust.
@@
Results. We show the results for the CS task in Figure 3 and provide an example posterior of RVNP with $N_{\rm obs}=1000$ in Figure 2. We display the simulated observations together with observed points in Figure 16. In this task, the misspecification of removing cells in the core regions that mimics necrosis primarily impacts the number of cancer cells summary statistics. In Figure 8, we show the evaluation of RVNP on a well-specified version of the CS task. In Figure 12, we show the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
All versions of our algorithm can recover robust posterior inference after the $N_{\rm obs}=10$ point in the CS task, performing better than NNPE and NPE. However, NNPE still performs relatively well, as the misspecification can be described by inflating the covariance along two specific axes. Both the global and the neural error covariance matrix represent a constant covariance matrix that points in a given direction that best describes the misspecification, which offers a form of model criticism for this problem. Our methods have better posterior coverage than NNPE and NPE, while at the same time achieving high log-probability and samples that are closer to the truth. Furthermore, they do not rely on any parameters that control the influence of the misspecification on the posterior.
Figure 3: Results for the CS task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling.
@@ 4.3 Task B: SIR
We include the misspecified susceptible-infected-recovered (SIR) task from ward2022robust, which takes the stochastic model of epidemic spread modelled conditional on $\bm{\theta}=(\beta,\gamma)$, the time-varying infection rate and the recovery rate, respectively. The SIR model emulates ideal disease transmission dynamics from the susceptible ($s$), infected ($i$), and recovered ($r$) parameters as
@@
ward2022robust employs a stochastic extension by using time-dependent transmission dynamics through a variable infection rate $\tilde{\beta}_{t}$, accounting for external factors such as policy interventions or pathogen mutations. This stochastic process is characterised using the basic reproduction number $R_{0t}=\frac{\tilde{\beta}_{t}}{\gamma}$, which follows the mean-reverting stochastic differential equation:
@@
where $\eta$ controls the mean reversion strength of $R_{0t}$ toward the equilibrium value $\frac{\beta}{\gamma}$, $\sigma$ represents the volatility parameter, and $W_{t}$ denotes standard Brownian motion. $\eta=0.05$ and $\sigma=0.05$ are fixed and the goal is to infer the parameters $\beta$ and $\gamma$. The Julia code to sample from this process was taken directly from the data products of ward2022robust. The summary statistics produced in this task are the mean, median and maximum number of infections, the day of the maximum number of infections, and the day at which half of the total number of infections was reached, and the mean autocorrelation of infections with lag 1.
Figure 4: Results for the SIR task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling, not to be confused with the name of the task.
@@
Results. This misspecification in this task only has a significant impact on the autocorrelation summary statistic. We show the results for the SIR task in Figure 4 and provide an example of the posterior distribution ($N_{\rm obs}=1000$) in Figure 2. We display the simulated observations together with the observed points in Figure 17. In Figure 9, we show the evaluation of RVNP on a well-specified version of the SIR task. In Figure 13, we plot the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
In this task, the misspecification appears extremely OOD along the axis of the autocorrelation statistic. All versions of our algorithm can recover robust posterior inference after the $N_{\rm obs}=10$ point, and they perform better than NNPE and NPE in this task. However, NNPE still performs well, as the misspecification can be described by inflating the covariance along the final axis. NPE performs very poorly in this task, due to the extreme misspecification along the final axis.
Our methods have better posterior coverage than NNPE and NPE, while at the same time achieving high log-probability, and samples that are closer to the truth. Furthermore, the global and neural error covariance matrix learn that the covariance inflates along the final axis. However, the neural error model outperforms the global error model, as the degree of misspecification along that axis depends on $\bm{\theta}$. The effect of this is that both the log-probability and the NRMSE are better for RVNP than the global version, as the global version scales up the covariance to account for this misspecification everywhere.
@@ 4.4 Task C: Pendulum Task
We describe a stochastic pendulum simulator that, given $\bm{\theta}:=[\omega_{0},A]$, samples the horizontal position of a frictionless pendulum at 200 time points evenly sampled every 0.05 seconds $\bm{x}_{\rm sim}=(f(t_{0}),...,f(t_{200}))$ where $f(t)=A\cos(\omega_{0}t+\phi)$ for $\phi\sim U(-\pi,\pi)$. In this task, $\omega_{0}$ and $A$ denote the fundamental frequency and amplitude, respectively, of the frictionless pendulum. $\phi$ is a stochastic phase shift. This task was inspired by the task from wehenkel2025addressing. However, it differs significantly and was adjusted to test our claim that increasing the number of posterior observations will better constrain the parameters.
@@
Misspecification We synthesise a time calibration error in the instrumentation that causes the instrument to take 200 measurements every 0.075 seconds instead of the simulated 0.05 seconds. There is a significant probability that the misspecified point will appear ID due to the form of the misspecification.
@@
Neural Statistic Estimation In this example, each of the data points is a single observation. The neural statistic estimator is an embedding of the full pendulum time series into a lower-dimensional representation of the data. Following chen2021neuralapproximatesufficientstatistics, we use the Shannon-Jensen InfoMax objective (hjelm2018learning) to target sufficient neural statistics, $\iota_{\bm{\omega}}$, to encode the high-dimensional data. This objective function maximises the mutual information between $\bm{x}_{\rm sim}$ and $\bm{\theta}$ using a discriminator network. We give an overview of the chosen architecture in the appendices.
@@
Figure 5: Results for the Pendulum task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling
Results. We show the results for the Pendulum task in Figure 5. In Figure 2, we display an example of the posterior distribution in the $N_{\rm obs}=1000$ case. We show the simulated observations together with the observed points in Figure 18. In Figure 10, we evaluate RVNP on a well-specified version of the Pendulum task. In Figure 14, we plot the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
@@
In this task, the misspecification appears more complex geometrically. Moreover, due to the nature of the misspecification, most of the points will appear ID relative to the simulated training points. All versions of our algorithm can recover more robust posterior inference beginning after the $N_{\rm obs}=10$ point than the NNPE and NPE. NNPE struggles significantly, most likely due to the spike-and-slab error model being unable to account for such a complex misspecification. However, RVNP and RVNP-G recover robust posterior inference for this task. In particular, RVNP is well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range in parameter space. Furthermore, RVNP is reasonably well calibrated. Due to the complexity of the error model, the tuned variant, RVNP-T, does not see the same degree of performance increase. However, it does perform better than both NPE and NNPE. NPE performs very poorly in the pendulum task. Both covariance models have their largest eigenvector pointing in the direction of changing fundamental frequency, as expected.